How much heat is needed to completely vapourize 3.6 g of propane (C3H8) at its boiling point?

Given: ΔHvap = 15.7 kJ/mol


1
Expert's answer
2021-04-02T05:11:21-0400

Molar mass of Propane gas =

3 ×12 + 1×8 = 44 g / mol.


Moles of Propane gas = 3.6 / 44

= 0.0818 mol.


Heat needed to Vapourise =

( 15.7 × 0.0818 ) Kj.


= 1.29 Kj .